OpenStudy (anonymous):

a bullet of mass 0.04 kg travelling at 300m/s hits a fixed wooden block and penetrates a distance of 4cm.find the average resistance of the wood

OpenStudy (anonymous):

Use the equation v^2=u^2+2as. Here v is zero, u is 300m/s and s is 4cm or 0.04m. Thus a=-300/0.04=-7500m/s^2. Now we can use F=ma to find the resistance, namely F=0.04*(-7500)=-300 N.

OpenStudy (anonymous):

Wait i made a mistake. I forgot to do u^2

OpenStudy (anonymous):

1/2*m*v^2=F*d 1/2*0.04*300^2=F*0.04 F=45000N

OpenStudy (anonymous):

Yeah, what the pizza guy said

OpenStudy (jamesj):

For the record, we're constantly discussing questions like this over in the Physics group. Feel free to post more like it over there.